# PLRD: Partially Linear Regression Discontinuity Inference

### v1

@@ 3 Relationship to Existing Methods
$(\widehat{a}_{w},\widehat{\beta}_{w})=\mathop{\rm argmin}_{(a,\beta)\in\mathbb{R}^{2}}\ \left\{\sum_{i\,:\,W_{i}=w}K\left(\frac{|X_{i}-c|}{h_{n}}\right)\left(Y_{i}-a-\beta(X_{i}-c)\right)^{2}\right\},\ \ \ \ \widehat{\tau}_{\texttt{LLR}}=\widehat{a}_{1}-\widehat{a}_{0},$ (7)
@@
where $h_{n}\rightarrow 0$ is a bandwidth that “localizes” the regression and $K(\cdot)$ is a kernel function with support on [1, 1]11[-1,\,1][ - 1 , 1 ]. In the decade following Hahn et al. (2001), local linear (or polynomial) regression remained the prevalent approach approach to inference in RDDs, with confidence intervals produced by applying off-the-shelf tools for heteroskedasticity-robust inference directly to the above regression (Imbens and Lemieux, 2008; Lee and Lemieux, 2010). A limitation of this approach, however, is that it leads to contradictory guidance on the bandwidth choice under a non-parametric specification as given in, e.g., Assumption 1. The accuracy of $\widehat{\tau}_{\texttt{LLR}}$ depends delicately on the bandwidth $h_{n}$ (Imbens and Kalyanaraman, 2012) and, unfortunately, for choices of $h_{n}$ that optimize this accuracy the bias and standard error of $\widehat{\tau}_{\texttt{LLR}}$ are of the same order—so off the shelf local linear regression inference will not have valid coverage for this choice of $h_{n}$. The practitioner is thus forced to invoke “undersmoothing”, whereby they voluntarily use a non-accuracy-optimizing choice of $h_{n}$ to justify validity of their inferential procedure.
@@
In a major advance, Calonico et al. (2014) and Armstrong and Kolesár (2018) introduced paradigms for inference in RDDs that formally account for the bias of local linear regression. As discussed in Stommes et al. (2023), these methods offer material improvements in the credibility of regression-discontinuity analyses over the earlier practice. The main idea in Calonico et al. (2014) is to estimate and correct for the bias of local linear regression by leveraging higher-order smoothness; we refer to this approach as the “bias-corrected” approach. Calonico, Cattaneo, Farrell, and Titiunik (2019) discuss the use of covariates under the bias-corrected approach to RDDs while Cattaneo and Titiunik (2022) provide a recent review. In contrast, the “bias-aware” approach developed by Armstrong and Kolesár (2018) involves conservatively widening intervals to account for worst-case bias (without needing to actually estimate the bias); see Kolesár and Rothe (2018), Imbens and Wager (2019) and Noack and Rothe (2024) for further applications of this paradigm.
@@
Interestingly, not only do the above papers under the bias-corrected vs. bias-aware paradigms explore different methodological strategies for inference in RDDs; they also focus on proving formal results under different formal paradigms. Calonico et al. (2014) and follow-ups focus entirely on asymptotic validity results—as we do here. They fix a data-generating process, and guarantee that they eventually get coverage to this data-generating process as the sample size gets large (as is also done in (6)). In contrast, most papers written under the bias-aware paradigm specify a quantitative smoothness class for the $\mu_{w}(x)$ (e.g., all functions with a second derivative bounded by a pre-specified constant $B$), and then seek guarantees that hold in finite samples and uniformly over this whole class. Getting uniform guarantees in finite samples of course requires stronger assumptions than getting pointwise asymptotic guarantees; and for this reason papers associated with the bias-aware paradigm often start by making stronger assumptions than those associated with the bias-corrected paradigm. We emphasize, however, that this is due to the sought formal guarantees—and it’s not that bias-aware methods inherently require stronger assumptions to work than bias-corrected methods.
@@
Our proposed method, PLRD, straddles these two literatures. Our confidence intervals are rooted in the bias-aware paradigm of Armstrong and Kolesár (2018), and we use numerical convex optimization to construct our estimator as in Imbens and Wager (2019). In contrast, our formal results are focused on pointwise asymptotics, and the assumptions and guarantee type given in our Theorem 1 closely mirror those in Theorem 1 of Calonico et al. (2014). One notable paper that also straddles these literatures in Armstrong and Kolesár (2020), who propose bias-aware confidence intervals for local linear regression with pointwise asymptotic guarantees. In particular, they show that if one runs local linear regression as in (7) with the mean-square-error optimal bandwidth choice $h_{n}$ following Imbens and Kalyanaraman (2012), then under generic conditions we can build valid, bias-aware 95% confidence intervals for $\tau_{c}$ as $\widehat{\tau}_{\texttt{LLR}}\pm 2.18\times\text{standard errors}$ rather than the conventional $\widehat{\tau}_{\texttt{LLR}}\pm 1.96\times\text{standard errors}$. We provide a comparison of PLRD to this simple rule of thumb in Appendix A.
@@
  • Because of this data-driven choice of $B$ in step (1), we are not able to provide uniform, finite-sample inferential guarantees; [^8] And, as shown in Armstrong and Kolesár (2018), this would in fact be impossible. thus, like in Calonico et al. (2014), we focus on pointwise asymptotic inferential statements.
@@
Regarding this last point, the earlier focus on the $M$-bounded curvature class appears to have been largely guided by intuitive simplicity of this class rather than any conceptual or empirical arguments. Given the empirical performance of PLRD on calibrated simulations, we suggest that our $B$-Lipschitz curvature class with linear treatment effects may be a good default choice in many applications. We conduct an ablation analysis to examine the effect of different choices on the behavior of our estimator in Appendix A.
@@ 4 Empirical Study
We now describe in more detail the experiment presented in the introduction (Table 1). As context for this experiment, Table 3 shows a replication of a number of high-profile RDD applications using a variety of methods; citations for each application are given in the leftmost column and further description of each benchmark dataset is provided Appendix B. We immediately see that the confidence intervals given by rdrobust, especially with the ${\it(b.c.)}$ interval choice, are much shorter than those given by RDHonest, thus naturally making the method attractive to practitioners. But how much can we trust these intervals?
Dataset rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
Lee (2008) $0.059\pm 0.022$ $0.059\pm 0.059$ $0.058\pm 0.032$ $0.079\pm 0.020$
Cattaneo et al. (2015) $7.507\pm 2.859$ $7.507\pm 3.413$ $8.030\pm 4.210$ $6.440\pm 2.374$
Matsudaira (2008) (read.) $0.030\pm 0.028$ $0.030\pm 0.032$ $0.031\pm 0.046$ $0.029\pm 0.033$
Matsudaira (2008) (math) $0.072\pm 0.025$ $0.072\pm 0.030$ $0.057\pm 0.041$ $0.077\pm 0.026$
Jacob and Lefgren (2004) (read.) $0.143\pm 0.042$ $0.143\pm 0.048$ $0.127\pm 0.046$ $0.093\pm 0.037$
Jacob and Lefgren (2004) (math) $0.203\pm 0.035$ $0.203\pm 0.039$ $0.187\pm 0.058$ $0.152\pm 0.032$
Oreopoulos (2006) $0.099\pm 0.104$ $0.099\pm 0.143$ $0.070\pm 0.125$ $0.072\pm 0.115$
Ludwig and Miller (2007) $-2.781\pm 2.363$ $-2.781\pm 2.682$ $-3.153\pm 2.828$ $-1.824\pm 2.019$
Meyersson (2014) $2.983\pm 2.797$ $2.983\pm 3.293$ $3.024\pm 3.109$ $3.162\pm 2.294$
@@
    Training: We train WGANs on each benchmark dataset to mimic the underlying data generating process. This involves generating running variables and outcomes for both treated and control units.
@@
The WGANs take as input random Gaussian noise $Z\in\mathbb{R}^{3}$. We train one generator $\widehat{g}_{x}(x)$ for the unconditional distribution of the running variable $X$, and two generators $\widehat{g}_{w}(x,z)$ ($w=0,1$) for the conditional distribution of the outcomes for treated and control units, given the running variable. We can calculate the ground truth $\widetilde{\tau}$ for the GAN estimated distribution by sampling random noise $Z_{q}\in\mathbb{R}^{3}$ for $q=1,\dots,Q$, where $Q$ is a large number ($10^{6}$ in our simulations), and generating outcomes at the threshold, as follows:
$\widetilde{\tau}:=Q^{-1}\sum_{q=1}^{Q}(\widehat{g}_{1}(0,Z_{q})-\widehat{g}_{0}(0,Z_{q})).$
Finally, we generate $i=1,\dots,N=10^{3}$ simulated datasets with confidence intervals $[a_{i,k},b_{i,k}]$, where the index $k$ stands for the $k$-th method we use to produce these CIs. The empirical coverage for the $k$-th method is given by $\widehat{p}_{k}=\frac{1}{N}\sum_{i=1}^{n}\mathbf{1}\{a_{i,k}\leq\widetilde{\tau}\leq b_{i,k}\}$. Table 1 shows the obtained results.
@@ 4.1 Revisiting existing simulation designs
For completeness, we also benchmark the PLRD estimator against other existing methods on the simulation settings originally used in Calonico et al. (2014, Section 6). [^9] Of these, Setting 1 was first used by Imbens and Kalyanaraman (2012), while Settings 2 and 3 were original to Calonico et al. (2014). In each of these settings, we draw the running variable $X_{i}\sim 2\,\text{Beta}(2,4)-1$ for $n=500$ samples, and the outcome is given by $Y_{i}=m(X_{i})+\varepsilon_{i}$, where $\varepsilon_{i}$ are mean-zero Gaussian noise with variance $0.1295^{2}$. The form of $m(\cdot)$ varies for the different simulation settings; see Calonico et al. (2014) for details on Settings 1 through 3. Setting 4 uses the function $m=m_{\text{quad}}$ as in Imbens and Kalyanaraman (2012, Section 4.6). Figure 2 plots one realization for each of these four different simulation settings.
Table 4 presents the empirical coverage and average width of 95% CIs produced by different methods across these simulation settings. As before, rdrobust falls short of nominal coverage; note that these numbers are in line with those originally reported by Calonico et al. (2014, Table 1). RDHonest delivers valid coverage at the expense of longer coverage. And, in Settings 1, 3 and 4, plrd delivers valid coverage with intervals whose length is competitive with rdrobust.
Our method, however, does not perform well in Setting 2. The reason for this poor performance is immediately revealed by Fig. 2: Our method relies on the linear treatment effect assumption (Assumption 2), whereas Setting 2 has strong curvature in the treatment effect that is manifested in a sharp jump in curvature at the cutoff. If one believes that such a sharp jump in the curvature change is realistic, one should not rely on Assumption 2; and in Appendix A we detail a variant of plrd without this assumption (and verify that it performs well in Setting 2 of Calonico et al. (2014)). Overall, however, we still rely on Assumption 2 in the default specification of plrd as we believe such sharp curvature effects in $\tau(x)$ to be unusual in most applications. [^10] The Setting 2 of Calonico et al. (2014) was originally derived by fitting a 5th-order polynomial to data from Ludwig and Miller (2007). However, unlike with our WGAN approach, the final simulation distribution as displayed in Figure 2(b) does not resemble the original data as summarized in Ludwig and Miller (2007, Figure IV(A)) in terms of, e.g., its signal-to-noise ratio. In the original dataset, the datapoints are counties; the running variable is poverty rate; the cutoff is associated with head-start eligibility; and the outcome is child mortality rate (per 100,000 children) from causes potentially associated with Head Start.
(a) Setting 1
(b) Setting 2
(c) Setting 3
(d) Setting 4
Figure 2: Visualization of four simulation settings from Calonico et al. (2014, Section 6) and Imbens and Kalyanaraman (2012, Section 4.6). In each setting, the conditional regression surface $m(\cdot)$ is displayed as a solid line, and the points represent data for one simulation realization.
rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
@@
Setting 1 86.3% 0.207 91.5% 0.244 95.2% 0.290 94.6% 0.235
Setting 2 91.1% 0.313 93.2% 0.342 96.9% 0.436 51.5% 0.309
Setting 3 88.6% 0.218 93.1% 0.250 95.7% 0.301 95.9% 0.239
Setting 4 87.8% 0.207 93.0% 0.246 96.8% 0.275 97.5% 0.236
Table 4: Empirical coverage and average width of 95% confidence intervals produced by different methods for the simulation settings in Calonico et al. (2014) and Imbens and Kalyanaraman (2012). Methods are run as in Table 1.
These simulation results highlight the strengths and limitations of our plrd method: It performs well in settings that align with its underlying assumptions and can be adapted to handle more complex scenarios. However, users should be cautious when applying plrd to settings with potential highly nonlinear treatment effects $\tau(x)$ (non-linearity in the baseline response function $\mu_{0}(x)$ is not a problem, as long as it is captured by the standard conditions required in Assumption 1).

### v2

@@ 3 Relationship to Existing Methods
$(\widehat{a}_{w},\widehat{\beta}_{w})=\mathop{\rm argmin}_{(a,\beta)\in\mathbb{R}^{2}}\ \left\{\sum_{W_{i}=w}K\left(\frac{|X_{i}-c|}{h_{n}}\right)\left(Y_{i}-a-\beta(X_{i}-c)\right)^{2}\right\},\ \ \ \ \widehat{\tau}_{\texttt{LLR}}=\widehat{a}_{1}-\widehat{a}_{0},$ (8)
@@
where $h_{n}\rightarrow 0$ is a bandwidth that “localizes” the regression and $K(\cdot)$ is a kernel function with support on [1, 1][-1,\,1]. In the decade following Hahn et al. (2001), local linear (or polynomial) regression remained the prevalent approach to inference in RDDs, with confidence intervals produced by applying off-the-shelf tools for heteroskedasticity-robust inference directly to the above regression (Imbens and Lemieux, 2008; Lee and Lemieux, 2010). A limitation of this approach, however, is that it leads to contradictory guidance on the bandwidth choice under a non-parametric specification as given in, e.g., Assumption 1. The accuracy of $\widehat{\tau}_{\texttt{LLR}}$ depends delicately on the bandwidth $h_{n}$, with a first systematic proposal given in Imbens and Kalyanaraman (2012). For choices of $h_{n}$ that optimize this accuracy the bias and standard error of $\widehat{\tau}_{\texttt{LLR}}$ are of the same order—so off-the-shelf local linear regression inference will not have valid coverage for this choice of $h_{n}$. The practitioner is thus forced to invoke “undersmoothing”, whereby they voluntarily use a non-accuracy-optimizing choice of $h_{n}$ to justify validity of their inferential procedure.
@@
In a major advance, Calonico et al. (2014) and Armstrong and Kolesár (2018) introduced paradigms for inference in RDDs that formally account for the bias of local linear regression. As discussed in Stommes et al. (2023), these methods offer material improvements in the credibility of regression-discontinuity analyses over the earlier practice. The main idea in Calonico et al. (2014) is to estimate and correct for the bias of local linear regression by leveraging higher-order smoothness; we refer to this approach as the “bias-corrected” approach. [^12] When reporting results using the bias-correct approach, we use robust standard error estimates as recommended by Calonico et al. (2014) and as implemented in the package rdrobust (Calonico et al., 2015; Calonico et al., 2017). Calonico, Cattaneo, Farrell, and Titiunik (2019) discuss the use of covariates under the bias-corrected approach to RDDs while Cattaneo and Titiunik (2022) provide a recent review. In contrast, the “bias-aware” approach developed by Armstrong and Kolesár (2018) involves conservatively widening intervals to account for worst-case bias (without needing to actually estimate the bias); see Kolesár and Rothe (2018), Imbens and Wager (2019) and Noack and Rothe (2024) for further applications of this paradigm.
@@
Interestingly, not only do the above papers under the bias-corrected vs. bias-aware paradigms explore different methodological strategies for inference in RDDs; they also focus on proving formal results under different formal paradigms. Calonico et al. (2014) and follow-ups focus entirely on asymptotic validity results—as we do here. They fix a data-generating process, and guarantee that they eventually get coverage for this data-generating process as the sample size gets large (as is also done in (7)). In contrast, most papers written under the bias-aware paradigm specify a quantitative smoothness class for the $\mu_{w}(x)$ (e.g., all functions with a second derivative bounded by a pre-specified constant $B$), and then seek guarantees that hold in finite samples and uniformly over this whole class. Getting uniform guarantees in finite samples of course requires stronger assumptions than getting pointwise asymptotic guarantees; and for this reason papers associated with the bias-aware paradigm often start by making stronger assumptions than those associated with the bias-corrected paradigm. We emphasize, however, that this is due to the sought formal guarantees; and it’s not that bias-aware methods inherently require stronger assumptions to work than bias-corrected methods.
@@
Our proposed method, PLRD, straddles these two literatures. Our confidence intervals are rooted in the bias-aware paradigm of Armstrong and Kolesár (2018), and we use numerical convex optimization to construct our estimator as in Imbens and Wager (2019). In contrast, our formal results are focused on pointwise asymptotics, and the assumptions and guarantee type given in our Theorem 1 mirror those in Theorem 1 of Calonico et al. (2014). One notable paper that also straddles these literatures in Armstrong and Kolesár (2020), who propose bias-aware confidence intervals for local linear regression with pointwise asymptotic guarantees. In particular, they show that if one runs local linear regression as in (8) with a mean-square-error optimal bandwidth choice $h_{n}$, then under generic conditions we can build valid, bias-aware 95% confidence intervals for $\tau_{c}$ as $\widehat{\tau}_{\texttt{LLR}}\pm 2.18\times\text{standard errors}$ rather than the conventional $\widehat{\tau}_{\texttt{LLR}}\pm 1.96\times\text{standard errors}$. We provide a comparison of PLRD to this approach in Appendix B.
@@
  • Because of this data-driven choice of $B$ in step (1), we are not able to provide uniform, finite-sample inferential guarantees; [^13] As shown in Armstrong and Kolesár (2018), doing so would in fact be impossible under smoothness assumptions like Assumption 1 alone. We do note that proving uniform bounds for methods involving data-driven smoothness guarantees can be possible under more restricted function classes, e.g., Armstrong and Kolesár (2020, Appendix E) consider such results over function classes where global polynomial approximations are guaranteed to capture worst-case local curvature of the underlying functions. We leave a study of uniformity guarantees for PLRD over restricted function classes to future work. thus, like in Calonico et al. (2014), we focus on pointwise asymptotic inferential statements.
@@
Regarding this last point, the earlier focus on the $M$-bounded curvature class appears to have been largely guided by intuitive simplicity of this class rather than any conceptual or empirical arguments. Given the empirical performance of PLRD on calibrated simulations, we suggest that our $B$-Lipschitz curvature class with linear treatment effects may be a good default choice in many applications. We conduct an ablation analysis to examine the effect of different choices on the behavior of our estimator in Appendix B.
@@ 4 Empirical Study
We now describe in more detail the experiment presented in the introduction (Table 1). As context for this experiment, Table 2 shows a replication of a number of high-profile RDD applications using a variety of methods; citations for each application are given in the leftmost column and further description of each benchmark dataset is provided Appendix C.
Dataset conventional rdrobust RDHonest plrd
Lee (2008) $064\pm 0.022 0.0$$59\pm 0.025 0.0$$58\pm 0.032 0.0$$073\pm 0.024$
Cattaneo et al. (2015) $7.414\pm 2.859$ $7.507\pm 3.413$ $8.030\pm 4.210$ $5.830\pm 2.127$
Ludwig and Miller (2007) $2.409\pm 2.363$ $2.781\pm 2.682$ $3.153\pm 2.828$ $1.249\pm 1.507$
Meyersson (2014) $3.020\pm 2.797$ $2.983\pm 3.293$ $3.024\pm 3.109$ $3.010\pm 2.093$
Matsudaira (2008) (read.) $036\pm 0.028 0.0$$30\pm 0.032 0.0$$31\pm 0.046 0.0$$017\pm 0.033$
Matsudaira (2008) (math) $075\pm 0.025 0.0$$72\pm 0.030 0.0$$57\pm 0.041 0.0$$073\pm 0.028$
Jacob and Lefgren (2004) (read.) $0.129\pm 0.042$ $0.143\pm 0.048$ $127\pm 0.046 0.0$$099\pm 0.040$
Jacob and Lefgren (2004) (math) $0.195\pm 0.035$ $0.203\pm 0.039$ $0.187\pm 0.058$ $0.168\pm 0.039$
Oreopoulos (2006) $070\pm 0.104 0.0$$99\pm 0.143 0.0$$70\pm 0.125 0.0$$021\pm 0.064$
@@
    Training: We train WGANs on each benchmark dataset to mimic the underlying data generating process for both treated and control units.
@@
Our WGAN training procedure consists of the following steps. First, apply marginal CDF transforms to the data to convert $(X_{i},Y_{i})$ to $(U_{i},V_{i})$, where $U_{i}=F_{x}(X_{i})$ and $V_{i}=F_{y}(Y_{i})$. For discrete data we break ties by randomization, and for additional stability we use a Gaussian transformation. We train two generators $\widehat{g}_{w}(u,z)$ ($w=0,1$), one for treated units and one for controls, that mimic the conditional distribution of the CDF-transformed outcome $V$ given the CDF-transformed running variable $U$. Finally, we invert the CDF-transforms to obtain $(\widetilde{X}_{i},\widetilde{Y}_{i})$ from the pair $(\widetilde{U}_{i},\widetilde{V}_{i})$ generated by the trained WGAN.
We calculate the ground truth $\widetilde{\tau}$ for the GAN estimated distribution as follows. We first calculate the cutoff $c^{\prime}$ under the applied CDF transforms. We then sample random noise $Z_{0,q},Z_{1,q}\in\mathbb{R}$ for $q=1,\dots,Q$, where $Q$ is a large number ($10^{7}$ in our simulations), generate CDF-transformed outcomes at the threshold $c^{\prime}$, and invert them:
$\widetilde{\tau}:=Q^{-1}\sum_{q=1}^{Q}\left(F_{y}^{-1}(\widehat{g}_{1}(c^{\prime},Z_{1,q}))-F_{y}^{-1}(\widehat{g}_{0}(c^{\prime},Z_{0,q}))\right).$
We generate $N=10^{4}$ simulated datasets indexed by $i=1,\dots,N$, and compute confidence intervals $a_{i,k},b_{i,k}]$, for each method $k$. The empirical coverage for the $k$-th method is given by $\widehat{p}_{k}=N^{-1}\sum_{i=1}^{n}\mathbf{1}\{a_{i,k}\leq\widetilde{\tau}\leq b_{i,k}\}$. Table 1 shows the obtained results.
@@ 4.1 Revisiting existing simulation designs
conventional rdrobust RDHonest plrd
@@
Pure noise 92.6% 1.419 93.1% 1.681 96.0% 1.854 97.2% 1.011
Setting 1 88.4% 0.203 90.9% 0.240 95.4% 0.288 94.8% 0.237
Setting 2 79.7% 0.266 90.5% 0.296 96.2% 0.440 97.3% 0.389
Setting 3 90.1% 0.213 92.7% 0.246 96.2% 0.301 99.6% 0.290
Setting 4 91.8% 0.201 92.3% 0.238 96.6% 0.276 96.1% 0.178
Table 3: Empirical coverage and average width of 95% confidence intervals produced by different methods on pure noise and on simulations from Calonico et al. (2014). All simulations have sample size $n=500$, and results are aggregated across 10,000 replications. Methods are run as in Table 1.
For completeness, we also benchmark the PLRD estimator against other existing methods on some standard simulation designs. First, we consider a pure noise setting where we generate the running variables as $X_{i}\sim\text{Unif}([-1,\,1])$, the outcomes as $Y_{i}\sim\mathcal{N}(0,\,1)$, and set treatment to $W_{i}=1(\{X_{i}\geq 0\})$. Next, we consider simulation settings used in Calonico et al. (2014, Section 6). In each of these settings, we draw the running variable $X_{i}\sim 2\,\text{Beta}(2,4)-1$, and the outcome is given by $Y_{i}=m(X_{i})+\varepsilon_{i}$, where $\varepsilon_{i}$ are mean-zero Gaussian noise with variance $0.1295^{2}$. The form of $m(\cdot)$ varies for the different simulation settings; see Calonico et al. (2014) for details on Settings 1 through 3. Setting 4 uses the function $m=m_{\text{quad}}$ as in Imbens and Kalyanaraman (2012, Section 4.6). We note that the linear treatment effect condition (Assumption 2) is not satisfied in several of these designs, so we expect our pragmatic safeguard (given in Algorithm 1) to play an important role in delivering robust performance.
Table 3 presents the empirical coverage and average width of 95% CIs produced by different methods across these simulation settings. As before, rdrobust falls short of nominal coverage; note that these numbers are in line with those originally reported by Calonico et al. (2014, Table 1). RDHonest delivers valid coverage at the expense of wider intervals, and plrd delivers valid coverage with intervals whose width is competitive with rdrobust.